Circular No. 7818

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

IM NORMAE A. Retter and S. J. O’Toole, University of Sydney; and S. Starrfield,

Arizona State University, report: “Medium-resolution optical spectra (400– 700 nm) were obtained of IM Nor on Jan. 26.75 UT using the Anglo-Australian Telescope (+ RGO spectrograph). Our spectra quantitatively resemble those reported by Duerbeck et al. (IAUC 7799). The strong hydrogen Balmer lines (Hα to Hδ) all show P-Cyg profiles, and we still see strong Fe II emission, confirming that this recurrent nova is a member of the Fe II class. This suggests that more material has been ejected than in a typical recurrent-nova outburst. The lack of evolution in the spectrum since those obtained by Duerbeck et al., in combination with the very slow decline in brightness, implies that it is a ‘slow’ recurrent nova and is retracing the outburst that occurred in 1920 (Elliot and Liller 1972, Ap.J. 175, L69). This would make it only the second slow nova among the class of recurrent novae (T Pyx being the other) and well worth studying. The light curve shown by Elliot and Liller indicates that the nova’s brightness remained near maximum for > 60 days. A 5000-s observation with the Chandra satellite is planned beginning on Feb. 4.75, and we strongly encourage observations at other wavelengths. Follow-up observations with Chandra will be scheduled if IM Nor is bright in x-rays.”

Broadband-V magnitudes from W. Liller, Viña del Mar, Chile: Jan. 20.312 UT, 8.75; 26.362, 8.13; 27.339, 8.60; 28.351, 8.42.

Visual magnitude estimates (cf. IAUC 7796): Jan. 21.803 UT, 8.7 (A. Pearce, Nedlands, W. Australia); 27.250, 9.2 (J. Garcia, Rama Caida, Argentina); 30.840, 9.5 (Pearce); 31.837, 9.4 (Pearce); Feb. 1.281, 9.3 (A. Amorim, Florianopolis, Brazil; moonlight); 2.309, 9.3 (Amorim); 2.834, 9.3 (Pearce); 3.340, 9.3 (Garcia); 3.832, 9.4 (Pearce).

SUPERNOVA 2002an IN NGC 2575 A. Gal-Yam and O. Shemmer, Tel Aviv University, write: “Full re-

duction of a spectrum (range 400–780 nm) of SN 2002an (cf. IAUC 7805), obtained with the Wise Observatory 1-m telescope (+ FOSC spectrograph) on Jan. 31 by J. Dann, shows a blue continuum. Correcting for the redshift of the host galaxy (z = 0.013; NED database), we also marginally detect a weak, narrow (FWHM ∼ 1500 km/sec) Hα line in emission. The spectrum is similar to spectra of the type-IIn supernova 1998S (e.g., Fassia et al. 2001, MNRAS 325, 907) around peak brightness.”

2002 February 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green